### v2

# OpenLLM-Ro - Technical Report on Open-source Romanian LLMs trained starting from Llama 2

Abstract
In recent years, Large Language Models (LLMs) have achieved almost human-like performance on various tasks. While some LLMs have been trained on multilingual data, most of the training data is in English. Hence, their performance in English greatly exceeds their performance in other languages. This document presents our approach to training and evaluating the first foundational and chat LLM specialized for Romanian.
1 Introduction
The Transformer architecture Vaswani et al. (2017) has become ubiquitous recently in the Natural Language Processing (NLP) domain, leading to state-of-the-art performance on various tasks, ranging from text classification to text generation. Decoder-only language models such as GPT series Radford et al. (2018, 2019); Brown et al. (2020) exhibit impressive capabilities such as understanding and generating natural language, learning and solving tasks they have not been directly trained on. In the last few years, Large Language Models (LLMs) development has exploded, with a plethora of models available Achiam et al. (2023); Chowdhery et al. (2023); Touvron et al. (2023); Hoffmann et al. (2022); Taori et al. (2023); Jiang et al. (2023).
Training such massive models requires a vast amount of available data (ranging from tens to hundreds of billions of text tokens) besides computational resources. Naturally, most work on such models focuses mainly on English, often excluding other languages. In this work, we focus on building the resources for building an LLM specialized in Romanian, a less-resourced language. By showcasing our approach and releasing our models, we strive to enable further research and exciting applications for Romanian-speaking users.
Our contributions can be summarized as follows:
  • We present and release111www.huggingface.co/OpenLLM-Ro/RoLlama2-7b-Base-v1 the first foundational Romanian LLM (i.e., RoLlama) based on the open-source Llama 2 model Touvron et al. (2023). Our collection of models is grouped under the OpenLLM-Ro initiative.
  • We collect and translate a large set of conversation and instructions, resulting in RoLlama2-7b-Instruct222www.huggingface.co/OpenLLM-Ro/RoLlama2-7b-Instruct-v1 and RoLlama2-7b-Chat333www.huggingface.co/OpenLLM-Ro/RoLlama2-7b-Chat-v1 models capable of following instructions and entertaining multi-turn conversations in Romanian.
  • We extend and adapt existing testing suites to evaluate the performance of different models on Romanian tasks.
  • We make our recipe444www.github.com/OpenLLM-Ro (i.e., datasets and code) and models publicly available to encourage further research both for Romanian and other low or less resourced languages.
2 Related Work
With the advent of open source LLMs, especially Llama2 Touvron et al. (2023) and Mistral Jiang et al. (2023), non-English specialized alternatives have been trained.  Luukkonen et al. (2023) have trained a family of Finnish foundational models, ranging from 186M to 176B parameters. Trained on a total of 38B tokens (with a custom tokenizer), their models showcase state-of-the-art performance on a variety of Finish downstream tasks. Similarly, Cui et al. (2023) have trained both a foundational and a chat variant of Llama2 for Chinese, while Wang (2023) worked on a domain-specific (i.e., media) Chinese LLM. Finetuned chat-models starting from Llama2 or Mistral have been developed for Norwegian555www.huggingface.co/norallm, Dutch666www.huggingface.co/collections/BramVanroy/geitje-7b-ultra-65c1ee010ad80fd1f6a8f208, Bulgarian777www.huggingface.co/INSAIT-Institute, and Serbian888www.huggingface.co/gordicaleksa/YugoGPT.
Unlike other language-specific work, we adopt an open-source approach and are also working on improving the foundational model (by employing continual pretraining) before building a chat variant.
3 Data
We employ two distinct data sources for training a Romanian LLM: documents for training the foundational model and instructions for the chat variant. The remainder of this section presents each dataset.
3.1 CulturaX
CulturaX Nguyen et al. (2023) represents a collection of documents in 167 languages. The authors combine mC4 Xue et al. (2021) and OSCAR Abadji et al. (2022) multilingual datasets using multiple cleaning and deduplication stages to ensure a high-quality corpus. The Romanian part of CulturaX consists of around 40M documents and 40B tokens.
3.2 Instructions and Conversations
We collect a large and diverse dataset of conversations and instructions to build the chat model. Since high-quality data in Romanian is very scarce, we resort to translating datasets such as Alpaca (both with GPT3.5999www.github.com/tatsu-lab/stanford_alpaca and GPT4101010www.github.com/Instruction-Tuning-with-GPT-4/GPT-4-LLM), Dolly111111www.huggingface.co/datasets/databricks/databricks-dolly-15k, NoRobots121212www.huggingface.co/datasets/HuggingFaceH4/no_robots, GPT-Instruct131313www.github.com/teknium1/GPTeacher, Orca141414www.huggingface.co/datasets/Open-Orca/SlimOrca-Dedup, Camel151515www.huggingface.co/datasets/camel-ai/math161616www.huggingface.co/datasets/camel-ai/chemistry171717www.huggingface.co/datasets/camel-ai/biology181818www.huggingface.co/datasets/camel-ai/physics, UltraChat191919www.huggingface.co/datasets/stingning/ultrachat/tree/main, and OpenAssistant Köpf et al. (2023)202020www.huggingface.co/datasets/OpenAssistant/oasst1. Additionally, we manually translated the MTBench corpora Zheng et al. (2023), leading to a total of around 2.3M instructions and conversations that were translated212121www.systransoft.com/ last accessed 6th of March 2024 and used for our experiments.
4 Training
Training the foundational and chat variants is done using the official Llama repository222222www.github.com/facebookresearch/llama-recipes. We pick a fixed learning rate from early experiments with a small portion of the data instead of the cosine scheduler proposed in the original paper. Due to hardware limitations, we experiment with a maximum sequence length of 512, 1024, and 2048, picking the biggest batch size that fits into the memory (i.e., 64 for the sequence length of 512, 32 for the sequence length of 1024, and 16 for the sequence length of 2048). We use the proposed AdamW optimizer, with weight decay set to 0.1 and gradient clipping. We do not employ LORA Hu et al. (2021) in any stage, as we perform full training for both the continual pre-training and supervised finetuning steps.
4.1 Foundational Model
Training the foundational model is done iteratively using continual pretraining on the CulturaX dataset. Starting from the base Llama2232323www.huggingface.co/meta-llama/Llama-2-7b-hf model, we train the first model on 5% of the CulturaX, the following model on the next 5% (leading to the second model seeing 10% of the data), and the third model on another 10% of the CulturaX. Therefore, we have three variants of a Romanian foundational model, variants that used 5%, 10%, and 20% of the entire CulturaX dataset. Training the foundational model is done over one epoch, with a fixed learning rate of 1e-4.
4.2 Instruct & Chat Models
We train RoLlama2-7b-Instruct using translated instructions while we further include the translated conversations from UltraChat and OpenAssistant to the finetuning dataset for RoLlama2-7b-Chat. Training is done with a fixed learning rate of 5e-5 over one epoch.
5 Evaluation
Evaluating a general LLM is not a straightforward task considering the vast number of benchmarks and evaluation protocols proposed to judge the quality of a model. Popular benchmarks include Open LLM Leaderboard242424www.huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard, an online leaderboard that encompasses six reasoning and general knowledge tasks: ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2020), TruthfulQA Lin et al. (2021), Winogrande Sakaguchi et al. (2021) and GSM8k Cobbe et al. (2021). MT-Bench Zheng et al. (2024) proposes a multi-turn question set and shows that strong LLMs (e.g., GPT-4) can act as judges and can approximate human preferences in a scalable and somewhat explainable way. AGIEval Zhong et al. (2023) is a benchmark designed to evaluate foundational models in the context of human-centric standardized exams. Evaluation is done in multiple settings, from zero and few-shot to Chain of Thought (COT) prompting. While benchmarks in the Open LLM Leaderboard select the most probable answer (out of four candidates) for each question, the focus on MT-Bench is more on the quality of the generated text than the intrinsic knowledge of the world of each model. Some experiments252525www.twitter.com/maximelabonne/status/1746317053143773628?s=19 try and compute a correlation factor between human preferences and results of previously mentioned benchmarks.
Evaluating LLMs in any language besides English only adds another layer of complexity. For Romanian versions of ARC, HellaSwag, MMLU, and TruthfulQA, we resort to ChatGPT translations Dac Lai et al. (2023), while we translate262626same system used for translating instructions Winogrande and GSM8k. The generally used few-shot values for ARC, HellaSwag, and MMLU (i.e., 25 shots for ARC, 10 shots for HellaSwag, and 5 shots for MMLU) lead to the prompt having a length equal to or even surpassing 2048 tokens. As some of our models have been trained only on sequences of length 512, 1024, or 2048, we opt to use multiple few-shot settings. More specifically, for ARC, we test with 0, 1, 3, 5, 10, and 25 shots, MMLU with 0, 1, 3, and 5, HellaSwag with 0, 1, 3, 5, and 10 shots, and Winogrande is evaluated in 0, 1, 3 and 5 shot setting. For GSM8k, we employ 1, 3, and 5 shot evaluation. Results in Tables 1, 2 are averaged over a few shot settings for each benchmark. More detailed results are presented in the Appendix.
Model Average ARC MMLU Wino HS GSM8k TQA
Existing Models
For Romanian
Okapi 36.05 36.88 31.17 56.28 48.91 0.43 42.64
Andrei481/Llama-2-7b-Romanian 36.53 35.49 30.86 55.98 47.99 0.83 48.01
Andrei481/Mistral-7B-v0.1-Romanian 37.42 37.80 33.34 56.37 49.42 0.99 46.63
General
Llama-2-7b 35.65 33.85 30.93 56.43 46.98 1.37 44.36
Llama-2-7b-chat 35.58 34.92 32.37 54.26 44.52 2.05 45.38
Mistral-7B-v0.1 42.66 41.22 40.84 60.20 53.63 13.82 46.24
Mistral-7B-Instruct-v0.2 45.63 43.09 44.87 59.26 54.12 10.86 61.56
Romanian LLMs
RoLlama2-7b-Base 38.32 35.83 30.47 60.16 55.52 2.17 45.78
RoLlama2-7b-Instruct 44.42 40.36 37.41 69.58 55.64 17.59 45.96
RoLlama2-7b-Chat 42.65 38.29 35.27 65.25 56.45 12.84 47.79
Table 1: Comparison between RoLlama and other LLMs on Romanian versions of downstream datasets (abbreviations: HS - HellaSwag, Wino - Winogrande, TQA - TruthfulQA). Bold denotes the best results within a category.
Model Seq. len % Data Average ARC MMLU Wino HS GSM8k TQA
RoLlama2-7b-Base 512 5 38.01 35.07 32.07 58.72 54.80 3.08 44.33
RoLlama2-7b-Base 1024 5 38.82 35.98 32.10 59.85 55.27 4.22 45.50
RoLlama2-7b-Base 2048 5 38.80 35.98 32.08 59.55 55.07 4.47 45.65
RoLlama2-7b-Base 1024 10 38.61 35.75 31.94 60.20 55.58 3.43 44.77
RoLlama2-7b-Base 1024 20 38.32 35.83 30.47 60.16 55.52 2.17 45.78
Table 2: Foundational models performance under various learning setups. Same notations as in Table 1.
Model Average 1st turn 2nd turn #Answers in Ro
Existing Models
For Romanian
Okapi 3.19 3.70 2.68 156 / 160
Andrei481/Llama-2-7b-Romanian 0.82 0.99 0.65 157 / 160
Andrei481/Mistral-7B-v0.1-Romanian 1.36 2.45 0.27 158 / 160
General
Llama-2-7b-chat 0.80 1.34 0.25 21 / 160
Llama-2-7b-chat (ro_prompted) 1.39 2.05 0.73 43 / 160
Mistral-7B-Instruct-v0.2 1.09 1.40 0.78 24 / 160
Mistral-7B-Instruct-v0.2 (ro_prompted) 5.84 6.06 5.63 151 / 160
Romanian LLMs
RoLlama-Instruct 4.31 5.66 2.95 160 / 160
RoLlama-Chat 3.91 4.25 3.57 154 / 160
Table 3: Chat models performance on translated MT-Bench in Romanian.
6 Results
Table 1 introduces the evaluation results for existing LLMs compared to the foundational and chat variants of RoLlama. Results show that training a Romanian LLM improves scores over general LLMs: around 3% for the foundational models and around 9% for the chat variant.
We observe a very high performance of Mistral models (both foundational and chat). However, since the authors of Mistral (Jiang et al. (2023)) do not provide many details about their training data, we cannot know if the model has been trained on Romanian data as well as other languages. Furthermore, it seems that finetuning Mistral on Romanian Alpaca 272727www.huggingface.co/Andrei481/Mistral-7B-v0.1-Romanian/discussions/1 drastically decreases the performance on considered benchmarks.
Turning to an evaluation based more on generation rather than just picking the most probable continuation, the performance on MT-Bench is presented in Table 3. We note that even when using a Romanian prompt, most Llama2 and Mistral completions or answers are either entirely in English or have the first few words in Romanian, then reverting to English. Specifically asking the models to answer in Romanian in the prompt somewhat alleviates this issue, especially for Mistral models. As seen in the last column in Table 3, when specifically prompted to answer in Romanian, Llama2 manages to generate text in Romanian in only 25% of cases, while Mistral answers are almost entirely in Romanian (over 90% of cases vs 70% when not prompted to answer in Romanian). We automatically detect the language of the answer282828www.pypi.org/project/langdetect/ and overwrite with 0 the score of each answers that is not in Romanian.
As expected, our models can mostly generate text in Romanian without being specifically told do so. The chat model also generates 6 texts in English, which might be explained by the presence of some translation errors in the training conversations. With the exception of the Mistral model, the second-turn scores for all models are much lower than the first-turn scores. For RoLlama2-7b-Instruct, this is somewhat expected as the model has predominantly seen single-turn conversations during training. While still underperforming, the first turn score of RoLlama2-7b-Instruct is within 0.5 points of Mistral. Adding multi-turn conversations (i.e., in the form of UltraChat and Open Assistant) increases the score on second turns, but the overall score is still lower. The same phenomenon can also be observed in Table 1. As the conversations account for about 70% of the finetuning data (around 1.6M samples), their quality directly affects the performance of the final model. Methods for validating and curating the translations and selecting high-quality samples for building the best data blend should be researched further as they could lead to lower training times and better-performing models.
7 Conclusions
This technical report presents initial experiments performed for building the first open-source LLM specialized for Romanian. While still in their infancy, the released models show promising results, outperforming existing solutions for various tasks. Pretraining on additional cleaner data, alignment with human preference Ziegler et al. (2019); Rafailov et al. (2024); Dong et al. (2023), finetuning with adapters Hu et al. (2021), validating existing translations or adding more instructions and conversations are just a few improvements that should increase the quality of RoLlama family of models. Note that we inherit the model and dataset licenses for all variants of RoLlama. Further work should include a more permissive combination of datasets for finetuning the Chat and Instruct variants.
Furthermore, we present a general recipe which we expect to work on different, more powerful LLMs as well (e.g., bigger Llama2, Mistral, Mixtral292929www.mistral.ai/news/mixtral-of-experts/ or Llama3303030ai.meta.com/blog/meta-llama-3/ models). We firmly believe that this is just the first step towards building and adapting LLMs for Romanian for both research and industry applications.
Acknowledgements
The universities contributed with researchers who worked pro-bono (POLITEHNICA Bucharest and the University of Bucharest) and computing power for training the models (POLITEHNICA Bucharest). Synchronization and collaboration took place within a project313131www.ilds.ro/llm-for-romanian carried out in the Institute of Logic and Data Science with funding from BRD Groupe Societe Generale.
We thank Alin Stefanescu and Laurentiu Leustean for participating in various discussions regarding this project.
Appendix A Detailed Results
In this section we present detailed results for Romanian versions of ARC-Challenge (Table 4), MMLU (Table 5), Winogrande (Table 6), HellaSwag (Table 7) and GSM8k (Table 8). ARC, HellaSwag, and Winogrande benchmarks are used to assess a model’s reasoning capabilities, as ARC contains science exam questions and HellaSwag and Winogrande consist of real-world scenarios framed as text completion and fill-in-a-blank tasks, respectively. MMLU benchmark covers 57 subjects ranging from STEM to social sciences, humanities, and others, with difficulty ranging from elementary to professional level testing both the problem-solving capabilities and world knowledge of a model. TruthfulQA is a benchmark devised to measure whether a model is truthful in answering questions, while GSM8k consists of school math problems and is used for the evaluation of multi-step mathematical reasoning. While ARC, HellaSwag, Winogrande, MMLU, and TruthfulQA are posed as multi-choice tasks (i.e., the model has to choose the most probable answer), GSM8k is framed as a generation task, in which the model has to generate the solution. Finally, the results represent in all tables the percentage of cases in which the model picks or generates the correct solution.
Model ARC-0 ARC-1 ARC-3 ARC-5 ARC-10 ARC-25 ARC Avg
Existing Models
For Romanian
Okapi 36.08 36.93 37.02 36.85 36.85 37.53 36.88
Andrei481/Llama-2-7b-Romanian 32.13 35.22 35.73 36.50 36.42 36.93 35.49
Andrei481/Mistral-7B-v0.1-Romanian 36.08 37.36 38.56 39.07 39.25 36.50 37.80
General
Llama-2-7b 31.36 33.93 34.10 34.36 34.45 34.88 33.85
Llama-2-7b-chat 32.73 35.22 34.53 35.90 35.13 35.99 34.92
Mistral-7B-v0.1 38.64 41.05 40.62 42.50 41.65 42.84 41.22
Mistral-7B-Instruct-v0.2 40.02 43.02 43.10 44.47 43.79 44.13 43.09
Romanian LLMs
RoLlama2-7b-Base 33.33 35.30 35.82 37.28 36.76 36.50 35.83
RoLlama2-7b-Instruct 40.27 39.16 40.36 40.27 41.30 40.79 40.36
RoLlama2-7b-Chat 36.85 38.22 38.39 38.39 38.99 38.90 38.29
Table 4: ARC performance under various few-shot settings.
Model MMLU-0 MMLU-1 MMLU-3 MMLU-5 MMLU Avg
Existing Models
For Romanian
Okapi 30.81 30.48 31.51 31.87 31.17
Andrei481/Llama-2-7b-Romanian 28.88 30.03 31.97 32.55 30.86
Andrei481/Mistral-7B-v0.1-Romanian 30.91 34.13 34.69 33.61 33.34
General
Llama-2-7b 27.93 30.62 32.16 32.99 30.93
Llama-2-7b-chat 30.47 31.53 33.41 34.06 32.37
Mistral-7B-v0.1 35.23 41.62 42.98 43.52 40.84
Mistral-7B-Instruct-v0.2 41.63 44.81 46.65 46.40 44.87
Romanian LLMs
RoLlama2-7b-Base 28.03 30.14 31.65 32.04 30.47
RoLlama2-7b-Instruct 33.82 37.70 38.70 39.42 37.41
RoLlama2-7b-Chat 30.54 35.64 37.41 37.47 35.27
Table 5: MMLU performance under various few-shot settings.
Model Wino-0 Wino-1 Wino-3 Wino-5 Wino Avg
Existing Models
For Romanian
Okapi 55.80 56.67 56.12 56.51 56.28
Andrei481/Llama-2-7b-Romanian 55.80 56.04 55.96 56.12 55.98
Andrei481/Mistral-7B-v0.1-Romanian 58.48 54.93 55.80 56.27 56.37
General
Llama-2-7b 55.09 57.06 57.22 56.35 56.43
Llama-2-7b-chat 53.75 55.01 54.06 54.22 54.26
Mistral-7B-v0.1 60.06 60.38 60.30 60.06 60.20
Mistral-7B-Instruct-v0.2 60.30 59.43 58.17 59.12 59.26
Romanian LLMs
RoLlama2-7b-Base 59.98 59.67 60.22 60.77 60.16
RoLlama2-7b-Instruct 70.96 69.38 68.90 69.06 69.58
RoLlama2-7b-Chat 65.59 65.11 66.14 64.17 65.25
Table 6: Winogrande performance under various few-shot settings.
Model HS-0 HS-1 HS-3 HS-5 HS-10 HS Avg
Existing Models
For Romanian
Okapi 49.11 48.31 49.06 48.87 49.20 48.91
Andrei481/Llama-2-7b-Romanian 47.76 47.58 48.11 48.24 48.27 47.99
Andrei481/Mistral-7B-v0.1-Romanian 51.56 48.27 49.51 48.99 48.77 49.42
General
Llama-2-7b 46.74 46.44 46.71 47.33 47.70 46.98
Llama-2-7b-chat 44.47 43.96 44.16 45.01 45.00 44.52
Mistral-7B-v0.1 52.72 53.02 53.69 54.25 54.49 53.63
Mistral-7B-Instruct-v0.2 54.44 53.60 53.96 54.07 54.51 54.12
Romanian LLMs
RoLlama2-7b-Base 54.98 55.11 55.80 55.74 55.97 55.52
RoLlama2-7b-Instruct 54.25 55.63 56.08 56.05 56.17 55.64
RoLlama2-7b-Chat 56.20 55.98 56.41 56.69 56.99 56.45
Table 7: HellaSwag performance under various few-shot settings.
Model GSM8k-1 GSM8k-3 GSM8k-5 GSM8k Avg
Existing Models
For Romanian
Okapi 0.68 0.15 0.45 0.43
Andrei481/Llama-2-7b-Romanian 0.76 0.91 0.83 0.83
Andrei481/Mistral-7B-v0.1-Romanian 0.99 1.29 0.68 0.99
General
Llama-2-7b 0.61 1.52 1.97 1.37
Llama-2-7b-chat 1.21 1.90 3.03 2.05
Mistral-7B-v0.1 6.14 17.06 18.27 13.82
Mistral-7B-Instruct-v0.2 0.45 14.40 17.74 10.86
Romanian LLMs
RoLlama2-7b-Vase 0.00 2.27 4.25 2.17
RoLlama2-7b-Instruct 12.66 19.56 20.55 17.59
RoLlama2-7b-Chat 9.40 14.03 15.09 12.84
Table 8: GSM8k performance under various few-shot settings.